# Pathwise uniform convergence of a full discretization for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise

### v4

@@ 3.2 A stability estimate for a discrete stochastic convolution
$\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\langle{g(1),g^{\prime}(\theta)}\rangle,$
where the second equality uses the linearity of $g(\theta)$. Hence, invoking Eq. 3.26 which implies $g(1)=e_{j+1}$, and using the definition of $g(\theta)$, we establish for any $\theta\in[0,1]$ that
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\big{(}I_{1}(\theta)+I_{2}(\theta)\big{)},$
where
$I_{1}(\theta)$ $:=\int_{t_{j}}^{t_{j+1}}\langle{e_{j+1},\,y_{h}^{3}(t_{j+1})-Y_{j+1}^{3}}\rangle\,\mathrm{d}t,$
$I_{2}(\theta)$ $:=\int_{t_{j}}^{t_{j+1}}\big{\langle}{e_{j+1},\,\Delta_{h}(Y_{j+1}-y_{h}(t))+Y_{j}-y_{h}(t)+y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\big{\rangle}.$
Given the inequality
$\langle e_{j+1},\,y_{h}^{3}(t_{j+1})-Y_{j+1}^{3}\rangle\leqslant 0,$
we infer that $I_{1}(\theta)\leqslant 0$. Using the definition of $\Delta_{h}$ and recognizing that $e_{j+1}=Y_{j+1}-y_{h}(t_{j+1})$, we derive the following for any $t\in[t_{j},t_{j+1}]$:
$\big{\langle}{e_{j+1},\,\Delta_{h}(Y_{j+1}-y_{h}(t))+Y_{j}-y_{h}(t)+y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\big{\rangle}$
@@
$\qquad{}+\big{\langle}{e_{j+1},\,Y_{j}-y_{h}(t)+y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\big{\rangle}.$
Employing Hölder’s inequality, Young’s inequality, and Poincaré’s inequality, we further deduce that
$\big{\langle}{e_{j+1},\,\Delta_{h}(Y_{j+1}-y_{h}(t))+Y_{j}-y_{h}(t)+y_{h}^{3}(t)-Y_{j+1}^{3}}\big{\rangle}$
@@
Subsequently, we establish that
$I_{2}(\theta)$ $\leqslant c\int_{t_{j}}^{t_{j+1}}\Big{[}\|y_{h}(t)-y_{h}(t_{j+1})\|_{\dot{H}_{h}^{1,2}}^{2}$
$\qquad\qquad+\|Y_{j}-y_{h}(t)\|_{L^{2}}^{2}+\|y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})\|_{L^{2}}^{2}\Big{]}\,\mathrm{d}t.$
By combining the above bounds for $I_{1}$ and $I_{2}$, we derive the differential inequality for $\|g(\theta)\|_{L^{2}}^{2m}$:
$\frac{\mathrm{d}}{\mathrm{d}\theta}\|g(\theta)\|_{L^{2}}^{2m}$ $\leqslant c\|g(\theta)\|_{L^{2}}^{2m-2}\int_{t_{j}}^{t_{j+1}}\Big{[}\|y_{h}(t)-y_{h}(t_{j+1})\|_{\dot{H}_{h}^{1,2}}^{2}$
$\qquad\qquad+\|Y_{j}-y_{h}(t)\|_{L^{2}}^{2}+\|y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})\|_{L^{2}}^{2}\Big{]}\,\mathrm{d}t.$
Applying Young’s inequality and using the definition $e_{j}=Y_{j}-y_{h}(t_{j})$, along with the specified $A_{j,m}$, $B_{j,m}$, and $C_{j,m}$, a straightforward computation yields the differential inequality:
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}$ $\leqslant c\tau\lVert{g(\theta)}\rVert_{L^{2}}^{2m}+c\tau\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\big{(}A_{j,m}+B_{j,m}+C_{j,m}\big{)}.$
Upon integrating this differential inequality over $[0,s]$ for any $s\in[0,1]$, we attain
$\lVert{g(s)}\rVert_{L^{2}}^{2m}\leqslant\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\int_{0}^{s}\left(\tau\lVert{g(\theta)}\rVert_{L^{2}}^{2m}+\tau\lVert{e_{j}}\rVert_{L^{2}}^{2m}+A_{j,m}+B_{j,m}+C_{j,m}\right)\,\mathrm{d}\theta.$
Employing Gronwall’s lemma subsequently gives
$\lVert{g(1)}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\tau\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\left(A_{j,m}+B_{j,m}+C_{j,m}\right).$
Taking the expectations on both sides yields
$\mathbb{E}\lVert{g(1)}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\tau\mathbb{E}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\left(A_{j,m}+B_{j,m}+C_{j,m}\right).$
The desired claim Eq. 3.27 then follows by recognizing that $g(1)=e_{j+1}$ and $g(0)=e_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)$.
Step 2. Fix any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$. For brevity, we denote $P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)$ as $\eta$. With the equality $\mathbb{E}[\lVert{e_{j}}\rVert_{L^{2}}^{2m-2}\big{\langle}{e_{j},\,\eta}\big{\rangle}]=0$, a direct calculation gives
$\mathbb{E}\left\lVert e_{j}+\eta\right\rVert_{L^{2}}^{2m}$ $=\mathbb{E}\left[\left\lVert e_{j}\right\rVert_{L^{2}}^{2}+2\langle e_{j},\eta\rangle+\left\lVert\eta\right\rVert_{L^{2}}^{2}\right]^{m}$
$=\mathbb{E}\left\lVert e_{j}\right\rVert_{L^{2}}^{2m}+\mathbb{E}\Big{[}\sum_{(\alpha,\beta,\gamma)\in\Lambda}C_{\alpha,\beta,\gamma}\left\lVert e_{j}\right\rVert_{L^{2}}^{2\alpha}\langle e_{j},\eta\rangle^{\beta}\left\lVert\eta\right\rVert_{L^{2}}^{2\gamma}\Big{]},$
where $\Lambda$ is the set of all triplets $(\alpha,\beta,\gamma)\in\mathbb{N}^{3}$ satisfying $\alpha+\beta+\gamma=m$ and $\alpha\neq m$, excluding the specific case $(m-1,1,0)$. Each $C_{\alpha,\beta,\gamma}$ is a positive integer constant dependent solely on the triple $(\alpha,\beta,\gamma)$ for each $(\alpha,\beta,\gamma)\in\Lambda$. Applying Cauchy’s inequality followed by Young’s inequality, and considering the constraint $-\frac{2\alpha+\beta}{\beta+2\gamma}\geqslant 1-m$ for all $(\alpha,\beta,\gamma)\in\Lambda$, we deduce that
$\mathbb{E}\bigg{[}\sum_{(\alpha,\beta,\gamma)\in\Lambda}C_{\alpha,\beta,\gamma}\lVert{e_{j}}\rVert_{L^{2}}^{2\alpha}\langle{e_{j},\eta}\rangle^{\beta}\lVert{\eta}\rVert_{L^{2}}^{2\gamma}\bigg{]}$ $\leqslant\mathbb{E}\bigg{[}\sum_{(\alpha,\beta,\gamma)\in\Lambda}C_{\alpha,\beta,\gamma}\lVert{e_{j}}\rVert_{L^{2}}^{2\alpha+\beta}\lVert{\eta}\rVert_{L^{2}}^{\beta+2\gamma}\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\sum_{(\alpha,\beta,\gamma)\in\Lambda}\Big{(}\tau\lVert{e_{j}}\rVert_{L^{2}}^{2m}+\tau^{-\frac{2\alpha+\beta}{\beta+2\gamma}}\lVert{\eta}\rVert_{L^{2}}^{2m}\Big{)}\bigg{]}$
$\leqslant c\tau\mathbb{E}\Big{[}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+\tau^{1-m}\lVert{\eta}\rVert_{L^{2}}^{2m}\Big{]}.$
Consequently, we arrive at the inequality
$\mathbb{E}\big{\lVert}{e_{j}+\eta}\big{\rVert}_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\tau^{1-m}\mathbb{E}\lVert{\eta}\rVert_{L^{2}}^{2m}.$
Considering the second term on the right-hand side of this inequality, we infer that
$\tau^{1-m}\mathbb{E}\lVert{\eta}\rVert_{L^{2}}^{2m}$ $\leqslant c\tau^{1-m}\mathbb{E}\Big{[}\int_{t_{j}}^{t_{j+1}}\lVert{F(Y_{j})-F(y_{h}(t))}\rVert_{\gamma(H,L^{2})}^{2}\,\mathrm{d}t\Big{]}^{m}$
$\leqslant c\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{F(Y_{j})-F(y_{h}(t))}\rVert_{\gamma(H,L^{2})}^{2m}\,\mathrm{d}t$
$\leqslant c\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)}\rVert_{L^{2}}^{2m}\,\mathrm{d}t$
$\leqslant c\tau\mathbb{E}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+A_{j,m},$
where the initial inequality is justified by Lemmas 3.1 and 3.3; the subsequent inequality holds due to Hölder’s inequality; the next one follows from Eq. 3.7; and the final inequality stems from the definition of $A_{j,m}$ in conjunction with the observation that $e_{j}=Y_{j}-y_{h}(t_{j})$. Combining the above estimates leads to the final inequality
$\mathbb{E}\big{\lVert}{e_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\big{\rVert}_{L^{2}}^{2m}$ (3.28)
[…]
Step 3. By combining Eq. 3.27 and Eq. 3.28, we obtain that for any positive integer $m$ and for all $0\leqslant j<J$,
$\|e_{j+1}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant(1+c\tau)\|e_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}+c\mathbb{E}(A_{j,m}+B_{j,m}+C_{j,m}).$
Applying the discrete Gronwall’s lemma and recalling that $e_{0}=0$, we deduce
$\max_{0\leqslant j\leqslant J}\|e_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant c\sum_{j=0}^{J-1}\mathbb{E}(A_{j,m}+B_{j,m}+C_{j,m}),\qquad\forall m\in\mathbb{N}_{>0}.$
@@
Taking into account Eqs. 3.19, 3.21 and 3.22, this establishes the validity of Eq. 3.23 for all $p=2m$, where $m$ is a positive integer. The generalization to the broader case $p\in[2,\infty)$ is achieved by invoking Hölder’s inequality. This completes the proof of the error estimate Eq. 3.23.
3.3.3 Proof of the stability estimate Eq. 3.24
@@
$Y_{j}=\eta_{j}^{(0)}+\eta_{j}^{(1)}+\eta_{j}^{(2)}\quad\text{for all }1\leqslant j\leqslant J.$
@@
Given the identity $\eta_{j}^{(0)}=(I-\tau\Delta_{h})^{-j}P_{h}v$ holding for all $0\leqslant j\leqslant J$, and using the stability property of $P_{h}$ from Lemma 3.1 along with Eq. 3.4 with $\alpha=\beta=0$, we assert that
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\eta_{j}^{(0)}}\rVert_{L^{q}}^{p}\text{ is uniformly bounded w.r.t.~{}$h$ and $\tau$ for all $p,q\in[2,\infty)$}.$ (3.29)
For any $p\in[2,\infty)$, from Lemma 3.1 and Eq. 3.5, it directly follows that
$\tau\sum_{j=0}^{J}\mathbb{E}\left[\lVert{Y_{j}-P_{h}Y_{j}^{3}}\rVert_{L^{2}}^{p}+\lVert{P_{h}F(Y_{j})}\rVert_{\gamma(H,L^{6})}^{p}\right]$ $\leqslant c\tau\sum_{j=0}^{J}\mathbb{E}\left[1+\lVert{Y_{j}}\rVert_{L^{6}}^{3p}\right].$
Moreover, one can establish that
$\tau\sum_{j=0}^{J}\mathbb{E}\lVert{Y_{j}}\rVert_{L^{6}}^{3p}$ $\leqslant c\max_{0\leqslant j\leqslant J}\lVert{Y_{j}}\rVert_{L^{3p}(\Omega;L^{6})}^{3p}$
$\leqslant c\max_{0\leqslant j\leqslant J}\left[\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{3p}(\Omega;L^{6})}^{3p}+\lVert{y_{h}(t_{j})}\rVert_{L^{3p}(\Omega;L^{6})}^{3p}\right]$
$\leqslant c\max_{0\leqslant j\leqslant J}\left[1+\lVert{y_{h}(t_{j})}\rVert_{L^{3p}(\Omega;L^{6})}^{3p}\right]$
$\leqslant c\left(1+\lVert{y_{h}}\rVert_{L^{3p}(\Omega;C([0,T];L^{6}))}^{3p}\right),$
where the third inequality is justified by Eq. 3.23, the condition $\tau\leqslant h^{2}$, and the application of the inverse inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant ch^{-1}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$. Consequently, by appealing to the regularity assertion in Eq. 3.18, we conclude that
$\tau\sum_{j=0}^{J}\mathbb{E}\left[\lVert{Y_{j}-P_{h}Y_{j}^{3}}\rVert_{L^{2}}^{p}+\lVert{P_{h}F(Y_{j})}\rVert_{\gamma(H,L^{6})}^{p}\right]$
remains uniformly bounded with respect to $h$ and $\tau$ for all $p\in[2,\infty)$. Using this result in conjunction with Lemmas 3.5 and 3.1, and taking advantage of the continuous embeddings of $\dot{H}_{h}^{2,2}$ and $\dot{H}_{h}^{1,6}$ into $L^{q}$ for any $q\in[2,\infty)$, we are able to assert that
$\tau\sum_{j=1}^{J}\mathbb{E}\left[\|\eta_{j}^{(1)}\|_{L^{q}}^{p}+\|\eta_{j}^{(2)}\|_{L^{q}}^{p}\right]$
is also uniformly bounded with respect to $h$ and $\tau$ for all $p,q\in[2,\infty)$. This finding, when coupled with the assertion in Eq. 3.29 and recognizing that $Y_{j}=\eta_{j}^{(0)}+\eta_{j}^{(1)}+\eta_{j}^{(2)}$ for $1\leqslant j\leqslant J$, permits us to affirm the stability estimate stated in Eq. 3.24.
@@
$Y_{j}-y_{h}(t_{j})=\xi_{j}^{(0)}+\xi_{j}^{(1)}+\xi_{j}^{(2)}.$ (3.30)
The sequences $(\xi_{j}^{(0)})_{j=0}^{J}$, $(\xi_{j}^{(1)})_{j=0}^{J}$, and $(\xi_{j}^{(2)})_{j=0}^{J}$ are recursively defined as follows:
$\begin{cases}\xi^{(0)}_{j+1}-\xi^{(0)}_{j}=\tau\Delta_{h}\xi^{(0)}_{j+1}+\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{(}y_{h}(t_{j+1})-y_{h}(t)\big{)}\,\mathrm{d}t,\quad 0\leqslant j<J,\\ \xi_{0}^{(0)}=0,\end{cases}$
@@
$\begin{cases}\xi^{(2)}_{j+1}-\xi^{(2)}_{j}=\tau\Delta_{h}\xi^{(2)}_{j+1}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant j<J,\\ \xi_{0}^{(2)}=0.\end{cases}$

### v5

@@ 4.2 A stability estimate for a discrete stochastic convolution
$\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\langle{g(1),g^{\prime}(\theta)}\rangle.$
Since $g(1)=\eta_{j+1}$, we have for any $\theta\in[0,1]$ that
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\biggl{[}\int_{t_{j}}^{t_{j+1}}\langle{\eta_{j+1},\,y_{h}^{3}(t_{j+1})-Y_{j+1}^{3}}\rangle\,\mathrm{d}t$
$+\int_{t_{j}}^{t_{j+1}}\big{\langle}{\eta_{j+1},\,\Delta_{h}\eta_{j+1}+Y_{j}-y_{h}(t)+y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\big{\rangle}\,\mathrm{d}t\biggr{]}.$
Using the fact $Y_{i}-y_{h}(t_{i})=\xi_{i}+\eta_{i}$ for all $0\leqslant i\leqslant J$, an elementary calculation yields
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\biggl{[}\int_{t_{j}}^{t_{j+1}}\langle{\xi_{j+1}^{2},\,y_{h}^{2}(t_{j+1})+Y_{j+1}^{2}}\rangle\,\mathrm{d}t$
$\qquad\quad{}+\int_{t_{j}}^{t_{j+1}}\lVert{\eta_{j}}\rVert_{L^{2}}^{2}+\lVert{\xi_{j}}\rVert_{L^{2}}^{2}+\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{2}}^{2}+\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{2}}^{2}\,\mathrm{d}t\biggr{]}.$
Hence, applying Young’s inequality and using the definitions of $A_{j,m}$, and $B_{j,m}$, we obtain the following differential inequality: for any $\theta\in[0,1]$,
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}$ $\leqslant c\tau\lVert{g(\theta)}\rVert_{L^{2}}^{2m}+c\tau\lVert{\eta_{j}}\rVert_{L^{2}}^{2m}+c\tau\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+c\big{(}A_{j,m}+B_{j,m}\big{)}.$
Employing Gronwall’s lemma subsequently gives
$\lVert{g(1)}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\tau\lVert{\eta_{j}}\rVert_{L^{2}}^{2m}+c\tau\lVert{\xi}\rVert_{L^{2}}^{2m}+c\left(A_{j,m}+B_{j,m}\right).$
Taking the expectations on both sides yields
$\mathbb{E}\lVert{g(1)}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\tau\mathbb{E}\lVert{\eta_{j}}\rVert_{L^{2}}^{2m}+c\tau\mathbb{E}\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\big{[}A_{j,m}+B_{j,m}\big{]}.$
The desired claim Eq. 4.28 then follows by recognizing that $g(1)=\eta_{j+1}$ and $g(0)=\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)$.
Step 2. Fix any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$. Using Hölder’s inequality, Young’s inequality, and the fact
$\mathbb{E}\biggl{[}\lVert{\eta_{j}}\rVert_{L^{2}}^{2m-2}\Big{\langle}{\eta_{j},\,\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rangle}\biggr{]}=0,$
a straightforward calculation yields
$\mathbb{E}\Big{\lVert}{\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}$
@@
For the second term on the right-hand side of this inequality, we can use the fact that $P_{h}$ is the $L^{2}$-orthogonal projection operator onto $X_{h}$. Additionally, we apply Lemma 4.3, Hölder’s inequality, and the Lipschitz continuity of $F$ from Eq. 4.7 to deduce that
$\tau^{1-m}\mathbb{E}\Big{\lVert}{P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}\leqslant c\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)}\rVert_{L^{2}}^{2m}\,\mathrm{d}t.$
Consequently,
$\mathbb{E}\Big{\lVert}{\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}$
@@
Using the fact $Y_{j}-y_{h}(t_{j})=\xi_{j}+\eta_{j}$, along with the definition of $B_{j,m}$, we readily conclude that, for any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$,
$\mathbb{E}\Big{\lVert}{\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}$ (4.29)
@@
Step 3. By combining Eq. 4.28 and Eq. 4.29, we arrive at the conclusion that for any $m\in\mathbb{N}_{>0}$ and for all $0\leqslant j<J$,
$\|\eta_{j+1}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant(1+c\tau)\|\eta_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}+c\tau\mathbb{E}\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\big{[}A_{j,m}+B_{j,m}\big{]}.$
Applying the discrete Gronwall’s lemma and recalling that $\eta_{0}=0$, we deduce that
$\max_{0\leqslant j\leqslant J}\|\eta_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant c\sum_{j=0}^{J-1}\mathbb{E}(\tau\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+A_{j,m}+B_{j,m}),\quad\forall m\in\mathbb{N}_{>0}.$
Taking into account the inequalities Eqs. 4.18, 4.21 and 4.27, we then obtain that
$\max_{0\leqslant j\leqslant J}\|\eta_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant c\tau^{m}+c\mathbb{E}\Big{[}\sum_{j=0}^{J-1}A_{j,m}\Big{]},\quad\forall m\in\mathbb{N}_{>0}.$ (4.30)
Step 4. Employing an analogous argument to that in Steps 1 and 2, and using the stability property of $P_{h}$ as stated in Lemma 4.1, as well as the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we can conclude that (see, also, [33, Lemma 4.1]) for any $m\in\mathbb{N}_{>0}$,
$\mathbb{E}\Bigl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}}\rVert_{L^{2}}^{2m}\Bigr{]}+\mathbb{E}\Bigl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}}\rVert_{L^{4}}^{4}\Bigr{]}$
is uniformly bounded with respect to $h$ and $\tau$. By Hölder’s inequality, we further infer that for any $p_{0}\in[2,\infty)$, there exists $q_{0}\in(2,\infty)$ such that
$\mathbb{E}\Bigl{[}\tau\sum_{j=1}^{J}\lVert{Y_{j}}\rVert_{L^{q_{0}}}^{p_{0}}\Bigr{]}$
is uniformly bounded with respect to $h$ and $\tau$. In view of this assertion and Eq. 4.27, we can use Hölder’s inequality to infer that
$\mathbb{E}\Big{[}\tau\sum_{j=0}^{J-1}\langle{\xi_{j+1}^{2},\,Y_{j+1}^{2}}\rangle^{m}\Big{]}\leqslant c\tau^{m},\quad\forall m\in\mathbb{N}_{>0}.$
Similarly, using the stability estimate of $y_{h}$ in Eq. 4.17, we have
$\mathbb{E}\Big{[}\tau\sum_{j=0}^{J-1}\langle{\xi_{j+1}^{2},\,y_{h}^{2}(t_{j+1})}\rangle^{m}\Big{]}\leqslant c\tau^{m},\quad\forall m\in\mathbb{N}_{>0}.$
Consequently,
$\mathbb{E}\Big{[}\sum_{j=0}^{J-1}A_{j,m}\Big{]}\leqslant c\tau^{m},\quad\forall m\in\mathbb{N}_{>0},$
which, combined with Eq. 4.30, further implies that
$\max_{0\leqslant j\leqslant J}\lVert{\eta_{j}}\rVert_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant c\tau^{m},\quad\forall m\in\mathbb{N}_{>0}.$
@@
Combining this inequality with Eq. 4.27, and noting that $Y_{j}-y_{h}(t_{j})=\xi_{j}+\eta_{j}$ for all $0\leqslant j\leqslant J$, we confirm the validity of Eq. 4.22 for all $p=2m$, where $m$ is a positive integer. The generalization to $p\in[2,\infty)$ is achieved by invoking Hölder’s inequality. This completes the proof of the error estimate Eq. 4.22.
4.3.2 Proof of the stability property Eq. 4.23
@@
$Y_{j}=\eta_{j}^{(0)}+\eta_{j}^{(1)}+\eta_{j}^{(2)}\quad\text{for all }0\leqslant j\leqslant J.$
@@
Given the identity $\eta_{j}^{(0)}=(I-\tau\Delta_{h})^{-j}P_{h}v$ for $0\leqslant j\leqslant J$, and using the stability property of $P_{h}$ from Lemma 4.1 along with the resolvent estimate Eq. 4.4 with $\alpha=\beta=0$, we assert that
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\eta_{j}^{(0)}}\rVert_{L^{q}}^{p}\biggr{]}\text{ is uniformly bounded w.r.t.~{}$h$ and $\tau$ for all $p,q\in(2,\infty)$}.$ (4.31)
For any $p\in(2,\infty)$, the stability property of $P_{h}$ from Lemma 4.1 and the growth property of $F$ in Eq. 4.5 imply that
$\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\left(\lVert{Y_{j}-P_{h}Y_{j}^{3}}\rVert_{L^{2}}^{p}+\lVert{P_{h}F(Y_{j})}\rVert_{\gamma(H,L^{6})}^{p}\right)\biggr{]}$ $\leqslant c+c\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}}\rVert_{L^{6}}^{3p}\biggr{]}$
Moreover, it can be established that
$\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}}\rVert_{L^{6}}^{3p}\biggr{]}$ $\leqslant c\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\left(\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{6}}^{3p}+\lVert{y_{h}(t_{j})}\rVert_{L^{6}}^{3p}\right)\biggr{]}$
$\leqslant c+c\lVert{y_{h}}\rVert_{C([0,T];L^{3p}(\Omega;L^{6}))}^{3p},$
where the second inequality is justified by the error estimate Eq. 4.22, the condition $\tau\leqslant h^{2}$, and the application of the inverse estimate $\lVert{u_{h}}\rVert_{L^{6}}\leqslant ch^{-1}\lVert{u_{h}}\rVert_{L^{2}}$ for all $u_{h}\in X_{h}$. Consequently, by appealing to the stability assertion Eq. 4.17, we conclude that
$\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\left(\lVert{Y_{j}-P_{h}Y_{j}^{3}}\rVert_{L^{2}}^{p}+\lVert{P_{h}F(Y_{j})}\rVert_{\gamma(H,L^{6})}^{p}\right)\biggr{]}$
remains uniformly bounded with respect to $h$ and $\tau$ for all $p\in(2,\infty)$. Using this result in conjunction with Lemmas 4.5 and 4.1, and taking advantage of the continuous embeddings of $\dot{H}_{h}^{2,2}$ and $\dot{H}_{h}^{1,6}$ into $L^{q}$ for any $q\in[2,\infty)$, we are able to assert that
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\left(\|\eta_{j}^{(1)}\|_{L^{q}}^{p}+\|\eta_{j}^{(2)}\|_{L^{q}}^{p}\right)\biggr{]}$
is also uniformly bounded with respect to $h$ and $\tau$ for all $p,q\in(2,\infty)$. Combining this finding with the assertion in (4.31) and noting that $Y_{j}=\eta_{j}^{(0)}+\eta_{j}^{(1)}+\eta_{j}^{(2)}$ for $1\leqslant j\leqslant J$, we confirm the stability property (4.23).
@@
$Y_{j}-y_{h}(t_{j})=\chi_{j}^{(0)}+\chi_{j}^{(1)}+\chi_{j}^{(2)}.$ (4.32)
The sequences $(\chi_{j}^{(0)})_{j=0}^{J}$, $(\chi_{j}^{(1)})_{j=0}^{J}$, and $(\chi_{j}^{(2)})_{j=0}^{J}$ are recursively defined as follows:
$\begin{cases}\chi^{(0)}_{j+1}-\chi^{(0)}_{j}=\tau\Delta_{h}\chi^{(0)}_{j+1}+\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{(}y_{h}(t_{j+1})-y_{h}(t)\big{)}+Y_{j}-y_{h}(t)\,\mathrm{d}t,\quad 0\leqslant j<J,\\ \chi_{0}^{(0)}=0,\end{cases}$
@@
$\begin{cases}\chi^{(2)}_{j+1}-\chi^{(2)}_{j}=\tau\Delta_{h}\chi^{(2)}_{j+1}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant j<J,\\ \chi_{0}^{(2)}=0.\end{cases}$